OpenStudy (anonymous):

How do I prove this identity? 1-sin(-t)/cos(-t)=cos(t)/1+sin(-t)

OpenStudy (anonymous):

sin (-t) = -sin(t) and cos(-t) = cos(t) but is it (1 - sin(-t))/cos(-t) = cos(t)/(1+sin(t))?

OpenStudy (anonymous):

Yes, that's what I have to prove.

OpenStudy (anonymous):

is it -t in the sin on the right?

OpenStudy (anonymous):

Wait, yeah, it is -t.

OpenStudy (anonymous):

okay... starting on the left... rewrite the sin and cos as indicated above (in my previous post). then multiply top and bottom by the conjugate of the numerator. it should work itself from there.

OpenStudy (anonymous):

Okay, I got it! Thank you!

OpenStudy (anonymous):

you're welcome